OpenStudy (anonymous):

Did I do this Algebra problem correctly? :) y=4(5^x-10) It's -36, right? :)

hartnn (hartnn):

y is -36 when x is 1

hartnn (hartnn):

sorry when x is 0

OpenStudy (anonymous):

Well, I got -40, but that wasn't part of the multiple choice responses. So... I got -36 because that's the closest. :P

hartnn (hartnn):

how did u get -40 ?? x^0=1 so 5^0=1 and 1-10=-9 4*-9=-36 .....

OpenStudy (anonymous):

y=4(5^x-10) Obviously 5 x 0 is 0 so we have 4 x -10 = 40.

hartnn (hartnn):

its 5^x not 5*x 5^x means 5*5*5*5....x times so if we do it 0 times , we get 1